# A Review of Barren Plateaus in Variational Quantum Computing

### v1

@@ VI Strategies to avoid or mitigate BPs
Ever since the discovery of BPs, there have been attempts to mitigate their detrimental effects. A common feature shared between the tools discussed here is that they break the assumptions which lead to the curse of dimensionality. First, we will review techniques which somehow transform the loss into the inner product between two objects that live in polynomially, rather than in exponentially, large spaces [11] (see Box VI). Second, we will review methods that mitigate BPs even when operating in exponentially large spaces. Here, we particularly refer the reader to the alternative initialization section, as this appear to be one of the most promising ways to jump start the training in a region of large gradients, rather than in a random place of the optimization landscape.
@@ VI.1 Shallow circuits
Perhaps one of the simplest ways to avoid BPs is to employ shallow PQCs. Here, even if the loss is concentrated for sufficiently deep versions of the circuit (either because it is too expressive or due to the effects of noise), if the number of gates is sufficiently small one can expect that that the effect of BPs is reduced for certain measurements. Examples of such cases are logarithmic-depth hardware-efficient ansatzes [46, 102, 56] with local measurements, or quantum convolutional neural networks [48, 60] which are shallow by design. Intuitively, we can understand that these circuits avoid BPs—even if randomly initialized—by constraining the space that the measurement operator can explore. These insights have been used in the literature to show that when using the variational quantum eigensolver, along with a shallow circuit, it is better to employ problem-encoding schemes that lead to more local measurement operators [47, 128].
It is important to note that while shallow circuits can provably have non-exponentially vanishing gradients, this is only true for certain measurements (for instance, a logarithmic-depth hardware efficient ansatz can still exhibit a BP for a non-local measurement [46]. The question still remains as to whether a solution exists within the set of parameters. By using few gates, it is entirely possible that there is no optimal set of parameters [83]. This phenomenon is known as a reachability deficit [21]. In general, there is no simple way of knowing whether the model will suffer from reachability deficits and a case-by-case analysis is necessary.
@@ VI.2 Small dynamical Lie algebras
While in the previous section we presented problem-inspired ansatzes which are still too expressive to avoid exponential concentration, there do exist cases of PQCs whose DLA’s dimension only grows polynomially with the number of qubits. In these cases, and critically depending on the choice of initial state and measurement operator, the circuit’s adjoint action can lead to small modules and thus avoid expressiveness induced BPs according to Eq. (13) even in its deeper forms.
One can attempt to reduce the circuit’s expressiveness by using circuits acting on smaller number of qubits [128, 129, 130], or encoding sufficient inductive biases about the problem one is trying to solve. In this context, it has been recently pointed out that one particularly important source of inductive biases is the set of symmetries that a problem satisfies. That is, if one knows that the task’s solution must respect some symmetry, then one should construct a variational models that only explore the symmetry-respecting solution space [131, 45, 132, 133, 134, 6, 135, 136]. For instance, it has been shown that permutation-equivariant unitary PQCs (circuits that remain invariant under qubit permutations) have a DLA whose dimension scales as $\mathcal{O}(n^{3})$ [40, 137]. Similarly, $U(1)$-equivariant circuits (i.e., circuits that preserve the Hamming weight of an initial state) have been shown to not exhibit BPs when they act on initial states whose Hamming weight does not scale with $n$[18, 138, 139]. To finish, we note that [140] studied a PQCs where all the gates commute, meaning that the size of the algebra is exactly equal to the number of distinct gate generators.
@@
Another archetypal example of a PQC with small DLA is that of parametrized matchgate circuits (also known as Ising-model HVAs [141, 142, 143, 144, 139, 145, 146]), where one can find that $\mathfrak{g}\simeq\mathfrak{so}(2n)$ and therefore $\dim(\mathfrak{g})\in\mathcal{O}(n^{2})$ [147, 70, 18, 39] (see Box VI.2). These circuits can be shown to have a reduced expressive power as they can only implement free-fermionic evolutions. Along a similar line, it has been shown that linear optics continuous variable PQCs acting on coherent light in $n$ modes have an associated algebra $\mathfrak{g}\simeq\mathfrak{o}(2n)$ [148, 149] thus also potentially leading to BP-free loss functions. Interestingly, these models have also been explored in classical machine learning, where they are known as efficient unitary neural networks [150, 151], and where it was already noted that they do not exhibit vanishing gradients at large depths.
While the previous examples showcase the fact that one can indeed find circuits with polynomial-sized DLAs, and thus to their action being reducible to some polynomial-sized space, it is worth noting that such cases are quite rare. In fact, as previously mentioned most circuits will exhibit exponentially large algebras [70, 18, 124], and the quest for other PQC architectures with small algebras is an open area of research.
@@ VI.3 Variable structure ansatzes
Variable structured PQCs attempt to mitigate expressiveness and noise induced BPs by leveraging classical machine learning protocol to iteratively grow the quantum circuit by placing (and removing) gates that empirically lower the loss function and keep large gradients during training. As such, these approaches are able to employ highly-expressive gate-sets by only exploring specific regions of the ansatz architecture hyperspace. Several machine learning-aided strategies have been employed within this context such as evolutionary algorithms, short-depth compilation algorithms, auto-machine learning, among others [152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164]. In particular, we highlight that one of the most-promising and well-studied variable ansatz strategy is the ADAPT-VQE [155, 154, 165, 166]. Here, one constructs a circuit for quantum chemistry applications with gates based on the UCCSD ansatz, and it has been shown that such strategy can lead to hardware-friendly circuits with few entangling gates which can be constrained to only explore regions of large gradients and potentially reach good solutions.
@@ VI.4 Alternative initialization strategies
Perhaps the greatest limitation of studying probabilistic BPs by analyzing the scaling of ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]$ or ${\rm Var}_{\bm{\theta}}[\partial_{\mu}\ell_{\bm{\theta}}(\rho,O)]$ is that these quantities can only tell us how concentrated the loss is in average across the whole landscape. Such information is of course important as it allows us to understand what to expect if we randomly initialize the model’s parameters. However, it is well known that randomly initializing the parameters is never a good idea, and that the initialization method can be critical in determining the model’s ultimate performance.
@@
Here we highlight that some of these initialization strategies have shown heuristic success and appear to be one of the most promising avenues to mitigate BPs. In fact, classical machine learning models can exhibit vanishing gradients issues just like their quantum counterpart (see below for more on this), and smart-initializations have been one of the main tools used therein. Hence, we expect that warm-start will play a fundamental role in the future of variational quantum computing. Still, care must be taken, as several limitations to these strategies have been reported [192, 193]. These include concerns such as the fact that one can initialize in a region which might have large gradients, but is not well connected to a global minima, or that might be classically simulable [11]. Indeed several workarounds have been proposed to prevent a smartly initialized circuit to run into problematic regions of the optimization landscape (such as monitoring the reduced state’s entanglement [28], or using state-of-the-art optimizers [194, 195, 196, 19]).
@@ VII Strategies that cannot avoid BPs
While many promising strategies have been devised to prevent or alleviate BPs, some intuitive approaches can fail due as they are unable to change the underlying reasons why a BP appears. Here we review two such strategies.
Changing the optimization method. It is common to find claims that it is possible to navigate a BP landscape (when randomly initialized) by changing the optimization method. For instance, one can assume that since the gradients are suppressed, one could find a loss minimizing direction by either using higher-order gradient information, gradient-free methods, quantum natural-gradients [197, 198] or other quantum fisher information-based schemes. However, it has been shown that such approaches can still require an exponential number of shots to obtain a loss minimizing direction [23, 199, 97].
Error mitigation and noise-induced BPs. Given the critical limitations that noise imposes on variational quantum algorithms, it is interesting to determine whether error mitigation techniques [200, 201, 202] are capable of sufficiently denoising the loss function to find a loss minimizing direction. Unfortunately, it has been shown that in the worse case an exponential number of resources need to be expended to mitigate the effect of noise-induced BPs [103, 203, 204, 205] even when the circuit is of sub-logarithmic depth for a class of worst-case circuits [206].
@@ VIII Exponential concentration elsewhere
The study of BPs has thus been informed by, and in turn impacted, work in other areas of quantum computing and machine learning. Here we briefly review some of these results.
@@
Quantum generative models. In quantum generative modeling one aims to optimize a quantum state such that its probability distribution (resulting from some set of measurements) matches a target distribution [207, 208, 209, 210, 211, 212, 213, 214, 54, 215, 216]. Quantum generative models are intrinsically implicit models as they only provide access to samples (and not the distributions themselves as do explicit models). Correspondingly, one can draw a distinction between explicit losses, which are formulated explicitly in terms of the model and target probabilities, and implicit losses, which compare samples from the model and the training distribution [212]. The tension between using an explicit loss, such as the Kullback-Leibler divergence, with an implicit model, was shown in [212, 90] for quantum circuit Born machines to lead to a BP. Conversely, implicit losses, such as the maximum mean discrepancy [210, 212] and some losses used for quantum generative adversarial networks [54, 90], can be viewed as the expectation value of an observable and so can be analyzed with the tools discussed previously.
@@
BPs have also been studied in quantum Boltzmann machines [215], where one uses trains a parametrized thermal quantum state model by minimizing a log-likelihood loss. In this case, the landscape is convex and so it is possible to derive convergence guarantees with only polynomial sample complexity [216]. Inherently quantum strategies for generative modelling also pose an interesting avenue for finding quantum generative models that avoid BPs [217, 212].
@@
Kernel-based quantum algorithms. Exponential concentration also arises in kernel-based schemes [218, 219, 220, 98, 221, 222, 223, 224, 225], where a quantum computer is used as an enhanced feature-space from which inner products or similarity measures between the data states can be estimated (a la kernel trick) [226, 227, 228]. Unlike PQCs, one strength of kernel methods is the optimal model is guaranteed to be found after training due to the convex training landscape. However, with exponentially concentrated kernels, this leads to a trivial model where the predictions on unseen inputs are independent of the training data inputs [219].
@@
Quantum optimal control. While exponentially concentrated optimization landscapes can arise in quantum optimal control schemes [229], their presence had not been truly studied until the BP literature started blossoming. The intrinsic connection between these two areas of study was first reported in [230] and later in [231]. Then, in [18] it was noted that tools from quantum optimal control, such as the study of the PQC’s DLA, can be used to diagnose BPs. Nowadays, the cross-fertilzation goes the other way and tools developed in variational quantum computing have been used in optimal control schemes [232, 233].
@@
Learnability. Insights from BPs have allowed researchers to present no-go theorems on the learnability of typical scrambling unitaries (i.e., typical random unitaries) [235, 53]. There is an interesting commonality between these results and statistical query learning [236] no-go theorems for learning typical random unitaries [237, 238, 239]. These no-go theorems can be viewed more exotically as having implications for the learnability of black hole dynamics [235, 240]. Here it is worth noting that while BPs are fundamentally an architecture dependent phenomenon (but can be applied to a wide range of problems), tools from statistical query learning can provide architecture-independent bounds for a given learning problem [239].
@@ IX Beyond barren plateaus in variational quantum computing
The study of BPs is clearly not the end of the story if one wants to guarantee that a variational quantum computing model can achieve a quantum advantage. Here we briefly review some of the other questions that need to be addressed.
@@
Local minima & solution quality. It has been shown that quantum landscapes can be plagued with exponentially many sub-optimal local minima, potentially making their optimization extremely hard [241, 17, 14, 242, 15, 243, 19]. Critically, the local minima issue appears for both shallow circuits where no BP arises but where no good solutions exist, but also in deep circuits where a solution might exist but reaching it is still computationally difficult. Indeed, we know that getting rid of the local minima by overparametrizing the PQC requires exponentially deep circuits [229, 83, 193, 244]. This realization actually leads to a very powerful insight: If we have a randomly initialized circuit with $\mathcal{O}(\operatorname{poly}(n))$ parameters which has a BP because the DLA is exponentially large (and thus the PQC is not overparametrized [83]), even if we allow for infinite measurements and get rid of the sampling noise, we will still not be able to train since the optimizer is exponentially likely to get stuck in a local minima [229, 83, 193, 244]. Addressing the problem of local minima will be fundamental to guarantee that a model will be useful (this will be particularly important when using alternative-initialization techniques).
@@
Classical simulability of BP-free models. As we have discussed above, many existing techniques that avoid the BP curse of dimensionality do so by encoding the relevant dynamics in some subspace that grows polynomially in the number of qubits $n$. Since an exponentially large space is no longer being used, one should be able to simulate the information processing ability of the PQC, and thus classically estimate the loss function [11, 245] (note that for non-classically simulable initial states or measurements the quantum computer may be needed to gather some data in an initial data collection phase). Not having to implement a PQC on the quantum device could lead to algorithms that are better suited for near-term quantum computers as the data acquisition phase could be less noisy than fully implementing the parametrized quantum circuit [11, 246, 247]. Elsewhere, recent results have showcased that classical algorithms, when equipped with measurements from a quantum computer, can be significantly more powerful than their fully classical counterparts [248, 218, 249, 250, 251].

### v2

@@ VI Strategies to avoid or mitigate BPs
Since the discovery of BPs, attempts have been made to mitigate their detrimental effects. A common feature shared between the tools discussed here is that they break the assumptions which lead to the curse of dimensionality. First, we will review techniques which somehow transform the loss into the inner product between two objects that live in polynomially, rather than in exponentially, large spaces [11] (see Box VI). While these approaches allow us to avoid BPs, we note that they also make variational models vulnerable to classical simulation (see more discussion in Sec. IX). Second, we will review methods that mitigate BPs even when operating in exponentially large spaces. Here, we particularly refer the reader to the alternative initialization section, as this appear to be one of the most promising ways to jump start the training in a region of large gradients, rather than in a random place of the optimization landscape.
@@ VI.1 Shallow circuits
Perhaps one of the simplest ways to avoid BPs is to employ shallow PQCs. Here, even if the loss is concentrated for sufficiently deep versions of the circuit (either because it is too expressive or due to the effects of noise), if the number of gates is sufficiently small one can expect that that the effect of both probabilistic and deterministic BPs is reduced for certain measurements. In addition, while this strategy does not help circumventing the deterministic BP arisen from the input state, the shallow depth PQCs can avoid the (unital) noise-induced BP. Examples include logarithmic-depth hardware-efficient ansatzes with local measurements [51, 29, 61], or quantum convolutional neural networks [53, 67] which are shallow by design (but also classically simulable [149]). Intuitively, these circuits avoid BPs—even if randomly initialized—by constraining the space that the measurement operator can explore. These insights have been used in the literature to show that when using the variational quantum eigensolver, along with a shallow circuit, using problem-encoding schemes that lead to more local measurement operators is more effective [52, 150].
It is important to note that while shallow circuits can provably have non-exponentially vanishing gradients, this is only true for certain measurements. For instance, a logarithmic-depth hardware efficient ansatz can still exhibit a BP for a non-local measurement [51]. The question still remains as to whether a good solution exists within the set of parameters. By using few gates, it is entirely possible that the set of parameters obtained from the optimization does not correspond to a good enough solution of the problem [91]. This phenomenon is known as a reachability deficit [132, 22]. In general, there is no simple way of knowing whether the model will suffer from reachability deficits and a case-by-case analysis is necessary. In addition, it has also been reported that shallow circuits can have spurious local minima [151, 91, 152] making it even harder to find any good solution that might, or not, exist in the landscape.
@@ VI.2 Small dynamical Lie algebras
In the previous section, we presented problem-inspired ansatzes that are still too expressive to avoid exponential concentration. However, there are cases of PQCs whose DLA dimensions grow only polynomially with the number of qubits. In these cases, depending critically on the choice of initial state and measurement operator, the circuit’s adjoint action may result in small modules. This allows the circuit to avoid probabilistic expressiveness-induced BPs, even in its deeper forms, as described by Eq. (13).
The circuit’s expressiveness can be reduced by using circuits with fewer qubits [150, 153, 154], or encoding sufficient inductive biases about the problem. In this context, it has been recently pointed out that one particularly important source of inductive biases is the set of symmetries that a problem satisfies. That is, if one knows that the task’s solution must respect some symmetry, then one should construct a variational models that only explore the symmetry-respecting solution space [155, 50, 156, 157, 158, 6, 159, 160]. For instance, it has been shown that permutation-equivariant unitary PQCs –circuits that remain invariant under qubit permutations– have a DLA whose dimension scales as $\mathcal{O}(n^{3})$ [45, 161]. Similarly, $U(1)$-equivariant circuits, have been shown to avoid BPs when the Hamming weight of the initial state does not scale with $n$ [18, 162, 163]. To finish, we note that [164] studied a PQCs where all the gates commute, meaning that the size of the algebra is exactly equal to the number of distinct gate generators.
@@
Another archetypal example of a PQC with small DLA is that of parametrized matchgate circuits (also known as Ising-model HVAs [165, 166, 167, 168, 163, 169, 170]), where one can find that $\mathfrak{g}\simeq\mathfrak{so}(2n)$ and therefore $\dim(\mathfrak{g})\in\mathcal{O}(n^{2})$ [171, 77, 18, 28] (see Box VI.2). These circuits can be shown to have a reduced expressive power as they can only implement free-fermionic evolutions. Along a similar line, it has been shown that linear optics continuous variable PQCs acting on coherent light in $n$ modes have an associated algebra $\mathfrak{g}\simeq\mathfrak{o}(2n)$ [172, 173] thus also potentially leading to BP-free loss functions. Interestingly, these models have also been explored in classical machine learning, where they are known as efficient unitary neural networks [174, 175], and where it was already noted that they do not exhibit vanishing gradients at large depths. We note, however, despite being able to avoid BPs, working with the polynomial DLA dimension could lead to the classical simulability of the VQAs (see Sec. IX).
While the previous examples illustrate that one can indeed find circuits with polynomial-sized DLAs, and thus to their action being reducible to some polynomial-sized space, it is worth noting that such cases are quite rare. In fact, as previously mentioned most circuits will exhibit exponentially large algebras [77, 18, 131], and the quest for other PQC architectures with small algebras is an open area of research.
@@ VI.3 Variable structure ansatzes
Variable structured PQCs attempt to mitigate probabilistic expressiveness and noise induced deterministic BPs by leveraging classical machine learning protocols to iteratively grow the quantum circuit by placing or removing gates that lower the loss function and keep large gradients. As such, these approaches are able to employ highly-expressive gate-sets by only exploring specific regions of the ansatz architecture hyperspace. Several machine learning-aided strategies have been applied, including evolutionary algorithms, short-depth compilation algorithms, and automachine learning [176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188]. In particular, we highlight that one of the most-promising and well-studied variable ansatz strategy is the ADAPT-VQE (and variants there-off for other circuits such as QAOA) [179, 178, 189, 190, 191, 192, 193]. For instance, one can construct a circuit for quantum chemistry applications with gates taken from some poot (e.g., from those arising in UCCSD), and it has been shown that this strategy has led to hardware-friendly circuits with few entangling gates that focus on regions with large gradients, often resulting in good solutions.
@@ VI.4 Alternative initialization strategies
One of the main limitations of studying probabilistic BPs by analyzing the scaling of ${\rm Var}_{\bm{\theta}}[\ell_{\bm{\theta}}(\rho,O)]$ is that these quantities only tell us how concentrated the loss is on average across the whole landscape. Such information is of course important as it allows us to understand what to expect if we randomly initialize the model’s parameters. However, it is well known that randomly initializing the parameters is never a good idea, and that the initialization method can be critical in determining the model’s ultimate performance.
@@
Here we highlight that some of these initialization strategies have shown heuristic success and appear to be one of the most promising avenues to mitigate BPs. In fact, classical machine learning models can exhibit vanishing gradients issues just like their quantum counterpart (see below for more on this), and smart-initializations have been one of the main tools used therein. Additionally, in chemistry problems such as finding a ground state of a molecular Hamiltonian, the initialization is done typically by taking into a problem structure such as initializing with a mean field approach. Hence, we expect that warm-start will play a fundamental role in the future of variational quantum computing. However, these strategies have limitations that must be considered [217, 218]. These include concerns such as the fact that one can initialize in a region which might have large gradients, but is not well connected to a global minima, or that might be classically simulable [11]. Indeed several workarounds have been proposed to prevent a smartly initialized circuit to run into problematic regions of the optimization landscape (such as monitoring the reduced state’s entanglement [34], or using state-of-the-art optimizers [219, 220, 19]).
@@ VII Strategies that cannot avoid BPs
While many promising strategies have been devised to prevent or alleviate BPs, some intuitive approaches fail because they do not address the underlying causes of BPs. Here we review two such strategies.
Changing the optimization method. It is common to find claims that it is possible to navigate a BP landscape (when randomly initialized) by changing the optimization method. For example, one might assume that despite suppressed gradients, it is possible to find a loss-minimizing direction by using higher-order gradient information, gradient-free methods, quantum natural-gradients [221, 222] or other quantum fisher information-based schemes. However, studies have shown that these approaches still require an exponential number of measurements to obtain a loss-minimizing direction [25, 223, 30].
Error mitigation and noise-induced BPs. Given the critical limitations that noise imposes on variational quantum algorithms, it is essential to investigate whether error mitigation techniques [224, 225, 226] can sufficiently denoise the loss function to find a loss-minimizing direction. Unfortunately, research has shown that in the worst cases, mitigating noise-induced BPs requires an exponential number of resources [109, 227, 228, 229]. This remains true even when the circuit has sub-logarithmic depth in certain worst-case scenarios [230]. Importantly, we note that it is not known whether such worst-case scenario is representative of average-cases.
@@ VIII Exponential concentration elsewhere
The study of BPs has both informed and been influenced by other areas of quantum computing and machine learning. Here we briefly review some of these results.
@@
Quantum generative models. In quantum generative modeling one aims to optimize a quantum state such that its probability distribution (resulting from some set of measurements) matches a target distribution [231, 232, 233, 234, 235, 236, 237, 238, 59, 239, 240, 241]. Quantum generative models are implicit because they only provide access to samples, unlike explicit models that provide direct access to the underlying distributions. Correspondingly, one can draw a distinction between explicit losses, which are formulated explicitly in terms of the model and target probabilities, and implicit losses, which compare samples from the model and the training distribution [236]. The tension between using an explicit loss, such as the Kullback-Leibler divergence, with an implicit model, was shown in [236, 98] for quantum circuit Born machines to lead to a BP. Conversely, implicit losses, such as the maximum mean discrepancy [234, 236] and some losses used for quantum generative adversarial networks [59, 98, 241], can be viewed as the expectation value of an observable and so can be analyzed with the tools discussed previously.
@@
BPs have also been studied in quantum Boltzmann machines [239], where a parametrized thermal quantum state model is trained by minimizing a log-likelihood loss. In this case, the landscape is convex and so it is possible to derive convergence guarantees with only polynomial sample complexity [240]. Inherently quantum strategies for generative modelling also pose an interesting avenue for finding quantum generative models that avoid BPs [242, 236].
@@
Kernel-based quantum algorithms. Exponential concentration also arises in kernel-based schemes [243, 244, 245, 105, 246, 247, 248, 249, 250], where a quantum computer is used as an enhanced feature-space from which inner products or similarity measures between the data states can be estimated (a la kernel trick) [251, 252, 253, 254]. Unlike PQCs, one strength of kernel methods is that for a given choice of a kernel function and dataset, one is guaranteed to find the optimal trained model that minimizes the empirical loss since the training landscape is convex. However, the kernel values can only ever be estimated statistically using a polynomial number of shots. If the kernel is exponentially concentrated, then the empirically estimated kernels contain, with high probability, no meaningful information about the input data. Training the model with these estimated kernel values then leads to a trivial data-insensitive model where the predictions are independent of the input data [244].
@@
Quantum optimal control. In quantum optimal control, exponentially concentrated optimization landscapes can arise [151], but their connection to BPs was only explored after the BP literature gained traction. The intrinsic connection between these two areas of study was first reported in [255] and later in [256]. Then, in [18] it was noted that tools from quantum optimal control, such as the study of the PQC’s DLA, can be used to diagnose BPs. Nowadays, some tools developed in variational quantum computing have started to see use in optimal control schemes [257, 258, 259, 256, 257, 260], and the extent of this cross-fertilization is yet to be seen.
@@
Learnability. Insights from BPs have allowed researchers to present theorems on the difficulty of variationally learning typical random states [51], typical random unitaries [262, 58], and the output distributions of random circuits [236]. There is an interesting commonality between these results and statistical query learning [263, 264] no-go theorems for learning typical random states [265], random unitaries [266, 267, 268], the expectation of an observable over random states [16] and the outputs distributions of random quantum circuits [269]. In all cases, it is worth noting that while BPs are fundamentally an algorithm-dependent phenomenon (but are, in some sense, independent of the learning problem), tools from statistical query learning can provide algorithm-independent bounds on the hardness of learning a specific function class. Nonetheless, mathematically there are close links between the two families of results, with the difficulty in both cases boiling down to a curse of dimensionality. For a thorough investigation of the link between barren plateaus and statistical query learning no-go theorems see Ref. [268].
@@ IX Beyond barren plateaus in variational quantum computing
The study of BPs does not fully address the challenges in ensuring that variational quantum computing models can achieve a quantum advantage. Here we briefly review some of the other questions that need to be addressed.
@@
Local minima & solution quality. It has been shown that quantum landscapes can be plagued with exponentially many sub-optimal local minima, potentially making their optimization extremely hard [270, 17, 14, 271, 15, 152, 272, 273, 19, 274, 273]. Critically, the local minima issue affects both shallow circuits, where no good solutions exist, and deep circuits, where a solution might exist but remains computationally difficult to reach. For example, when learning an unknown unitary with a shallow one-dimensional parametrized circuit, there are exponentially many sub-optimal local minimum [272]. Indeed, we know that if the PQC is too shallow, then spurious local minima can appear in the landscape as an effect of not having enough parameters to control all available directions. While these minima can be be removed by overparametrizing the PQC, i.e., by adding more parameters, this could requires exponentially deep circuits [151, 91, 218, 275, 274]. Such realization actually leads to a very powerful insight: If we have a randomly initialized circuit with $\mathcal{O}(\operatorname{poly}(n))$ parameters which has a BP because the DLA is exponentially large (and thus the PQC is not overparametrized [91]), even if we allow for infinite measurements and get rid of the sampling noise, we will still not be able to train since the optimizer is exponentially likely to get stuck in a local minima [151, 91, 218, 275, 273]. Addressing the problem of local minima will be fundamental to guarantee that a model will be useful (this will be particularly important when using alternative-initialization techniques).
@@
Classical simulability of BP-free models. As we have discussed above, many existing techniques that avoid the BP curse of dimensionality do so by encoding the relevant dynamics in some subspace that grows polynomially in the number of qubits $n$. Since an exponentially large space is no longer being used, one should be able to simulate the information processing ability of the PQC, and thus classically estimate the loss function [11] (note that for non-classically simulable initial states or measurements the quantum computer may be needed to gather some data in an initial data collection phase). This is possible for all parameter settings for shallow hardware efficient ansätze [276, 277] and problem inspired ansätze with polynomial DLAs [94, 278, 279, 207, 280], and possible for any randomly chosen parameter setting with high probability for a wide class of unstructured parameterized quantum circuits [281, 282], QCNNs [149], noisy circuits [283, 112] and small-angle initialization strategies [284].
Not having to implement a PQC on the quantum device could lead to algorithms that are better suited for near-term quantum computers as the data acquisition phase could be less noisy than fully implementing the parametrized quantum circuit [11, 285, 286]. Moreover, the connection between classical simulability ans trainability is a current area of analysis [287]. Elsewhere, recent results have showcased that classical algorithms, when equipped with measurements from a quantum computer, can be significantly more powerful than their fully classical counterparts when solving certain contrived tasks [288, 243, 289, 290, 291]. Still, while it remains to be determined if such separation holds for practical applications, this is regardless a promising direction of research.
Finally, it is worth highlighting that there are numerous fully classical machine learning approaches that tackle the same applications as variational quantum computing. These include variational tensor network methods [292, 293] and neural network states [294, 295, 296, 297] for solving ground state problems, simulating the dynamics of quantum systems and state tomography. While limited by the intrinsic challenges of classically simulating quantum systems, these approaches naturally sidestep the barren plateau problem completely.